The perimeter of a rectangle is 36 cm. The length of the rectangle is 3 cm more than twice the width. What are the dimensions of the rectangle?
Answer by bluemockingjay7(41)

perimeter = 36cm
length = 2x + 3
width = x
2x + 3 + x = 36
2(2x + 3 + x) = 36
4x + 6 + 2x = 36
6x = 30
6x/6 = 30/6
x = 5
width = 5 cm
length = 2(5) + 3 = 13 cm